Garcia, Inc. uses a job-order costing system for its products, which pass from the Machining Department, to the Assembly Department, to finished-goods inventory. The Machining Department is heavily automated; in contrast, the Assembly Department performs a number of manual-assembly activities. The company applies manufacturing overhead using machine hours in the Machining Department and direct-labor cost in the Assembly Department. The following information relates to the year just ended:

Machining Department Assembly Department
Budgeted manufacturing overhead $ 4,000,000 $ 3,080,000
Actual manufacturing overhead 4,280,000 3,050,000
Budgeted direct-labor cost (based on practical capacity) 1,500,000 5,600,000
Actual direct-labor cost 1,450,000 5,780,000
Budgeted machine hours (based on practical capacity) 400,000 100,000
Actual machine hours 425,000 110,000


The data that follow pertain to job no. 775, the only job in production at year-end.

Machining Department Assembly Department
Direct material $ 23,500 $ 6,600
Direct labor $ 27,800 $ 58,600
Machine hours 360 150


Selling and administrative expense amounted to $2,500,000.

1. Compute the cost of the company’s year-end work-in-process inventory.

2. Determine whether overhead was under- or overapplied during the year in the Machining Department.

3. Determine whether overhead was under- or overapplied during the year in the Assembly Department.

4. How much overhead would have been charged to the company’s Work-in-Process account during the year?
Answer #1
1
Machining Assembly
a Budgeted Manufacturing Overhead 4,000,000 3,080,000
b Budgeted Machine Hours 400,000
c Budgeted Direct labor cost           5,600,000
d Overhead Rate of
Machining =(a/b)
Assembly= (a/c)
10                      0.55
2
Machining Assembly Total
a Direct Material 23,500 6,600 30,100
b Direct Labor 27,800 58600 86,400
c Overhead cost for Machining 3600 3,600
d Overhead cost for Assembly           32,230.00 32,230
e Total cost of work in Process 54,900           97,430.00 152,330
Total cost of work in Process 152,330
Working note
overhead cost of Machining = (Machine hour * Overhead rate)
Machining= 360*10= 3600
Overhead cost of Assembly = (direct labor cost * overhead rate)
Overhead cost of Assembly = 58600*0.55=32230
3 & 4
Machining Assembly
a applied overhead cost
Machining = Overhead rate * actual Machine hours
= 10*425,000
Assembly = overhead rate * actual direct labor cost = 0.55*5780000
          4,250,000.00    3,179,000.00
b actual overhead cost           4,280,000.00     3,050,000.00
c Over or under applied overhead (a-b)                  (30,000.00)         129,000.00
3 Matching department under -applied because Budgeted Manufacturing overhead cost is lower than actual overhead cost
4 Assembly Department over- applied Budgeted Manufacturing overhead cost is higher than actual overhead cost
5 Overhead Charged to Work in process during the year
Machining Assembly Total
applied Manufacturing Overhead           4,250,000.00    3,179,000.00    7,429,000.00
Overhead Charged to Work in process during the year =7,429,000 (sum of Both department)